We are given the lengths of the 3 altitudes of a trgle ABC


we have to find any one side


Any Hints???

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Palash Chordia,

Ans:- The area of a triangle is Δ and let the altitudes are p,q,s and the sides are a,b,c

Hence 1/2 a p=1/2 b q =1/2 r c=Δ

or ap=bq=cr=Δ=√s(s-a)(s-b)(s-c)

where s=(a+b+c)/2

From the above equations any one side may be determined solvig some linear or quadratic equations.

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Palash Chordia !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur